OpenStudy (anonymous):

You buy a plastic dart gun, and being a clever physics student you decide to do a quick calculation to find its maximum horizontal range. You shoot the gun straight up, and it takes 4.1 s for the dart to land back at the barrel. What is the maximum horizontal range of your gun?

OpenStudy (anonymous):

Have you attempted to use the motion formulas? You might notice that: To find max. hor. distance, you'll need velocity and time. To find the velocity, you'll need time and acceleration. I'll give a short example on finding the initial velocity of the bullet just exiting the gun barrel. by using \[v = u+at\] where v=0 at the top, a=-10(gravitational acceleration is different in direction than the initial veloicty), t is given, \[0=u-10(4.1)\] \[u=41ms^{-1}\] A similar case follows the horizontal distance.